### 18 Dec 2023 v1

@@ 2.4.2 Action recognition models (spatio-temporal-based HAR)
Transformer-based methods: As previously noted, 3D CNN approaches have several drawbacks that limit their applicability when dealing with complex actions of variable durations. Transformer-based approaches bridge this gap by providing a unique solution. The Transformer model can process sequential data in parallel by introducing positional embeddings. An early transformer-based method [77], Video Action Transformer Network (VATN), consists of two main components: the Spatial Transformer and the Temporal Transformer. The learned transformers allow the network to selectively attend to important features, resulting in improved action recognition accuracy. Recently, the MTV [223] framework combines multiple views of a video sequence to capture comprehensive spatial and temporal information. This introduces a fusion mechanism that combines the outputs of the both transformers. Additionally,  [68] introduces a framework that utilizes masked auto-encoders to capture spatial and temporal dependencies in video data. This approach offers a valuable mechanism for leveraging masked auto-encoders with video data to enhance the understanding and recognition of intricate video patterns. We refer keen readers to [84] and [200] for a more detailed analysis of action recognition methods.
@@ 3.1 Applications and model frameworks
Initial studies have primarily employed hand-designed features based on motion-strength, motion-trajectory, or average differential luminance signals in video segments to develop automated seizure detection systems. For example, hand-crafted features are extracted based on patient motion trajectories by attaching infrared reflective markers to specific body key points and by exploiting RGB-D sensors (e.g. the Microsoft Kinect camera) to distinguish Temporal Lobe Epilepsy (TLE) and Extratemporal Lobe Epilepsy (ETLE) [50]. Aghaei et al. [3] proposed a multimodal seizure detection approach to detect nocturnal convulsive seizures by combining hand-crafted features from video and EEG signals, achieving a higher accuracy than using either modality alone. Similarly, spatiotemporal interest points are extracted from video sequences to construct a histogram of word frequency features for video representation to detect Rolandic epilepsy in [219]. Other works have focused on nocturnal convulsive seizures using hand-crafted detection algorithms [75, 214] by exploiting optical flow calculations and the reconstruction vector field of velocities [102]. The algorithm quantifies the oscillatory movements seen as vibrations during the tonic phase, and clonic movements in the clonic phase. Recently, Garção et al. [73] proposed a privacy-preserving video-based seizure-detection and registration method to classify generalized tonic–clonic and focal to bilateral tonic–clonic seizures. This approach is based on optical flow, dimensionality reduction with principal component analysis, seizure movement isolation with independent component analysis, and machine learning classification. However, these approaches are limited in performance due to their inability to generalize to changing clinical environments (e.g. changes to subject orientation and distance, dealing with multiple individuals in the field of view, accounting for lighting and furniture changes, motion blur, variable frame rate) and their reliance on handcrafted features.
@@
For early detection, Fang et al. [64] proposed a spatial temporal gated recurrent unit convolutional neural network capable of not only early detection, but also of localizing the onset time of epileptic seizures. However, the specific seizure types considered and whether the system detects the full expression of the semiology or just the seizure onset remain unclear. In a recent study, Mehta et al. [141] demonstrated the feasibility of detecting tonic-clonic seizures (TCS) during their progression based on optical flow data. The approach utilized knowledge distillation and transformed-based approaches, showcasing a novel direction for early detection while addressing privacy concerns.
@@
The creation of devices capable of objectively recording seizure counts and characteristics holds immense value, not only for seizures occurring during sleep but also for those manifesting while awake. There has been an increasing interest in the analysis of extensive video data collected at home, particularly for understanding seizure semiology in patients with multiple indicators [13]. A notable example is the Nelli®hybrid system [22], a semi-automatic seizure monitoring platform utilizing audio/video inputs. This hybrid system employs computer vision and machine learning to identify kinematic data associated with seizures. Human experts then visually assess these identified events. The utility for reviewing nocturnal video recordings and motor seizures has been demonstrated [16, 165, 119]. While Nelli demonstrates proficiency in detecting subtle motor seizures, a study noted a lower accuracy in classifying more discrete motor seizures such as myoclonic jerks, short tonic seizures, and epileptic spasms [158]. The system has received endorsement for clinical use in Finland from a government-established body, the National Coordinating Group for Drug-resistant Epilepsy.
@@
Seizure semiology classification: Understanding the progression of automated pipelines within the context of video analysis for seizure event classification is crucial. While action recognition requires extracting spatio-temporal features of movements for the classification of epilepsy types, some works overlook the temporal information within seizures, treating the task as an image classification problem. For example, Maia et al. [138] used an InceptionV3 feature extractor to extract features for all frames, which were subsequently concatenated and fed to a multilayer perceptron, to distinguish between TLE and ETLE within IR seizure videos. Similarly, Pothula et al. [164] differentiated between Gelastic and Dacrystic facial seizures using a feature extraction layer that recognizes the patient’s emotion through a facial emotion recognition model and classification layer (Random Forest) that received concatenated features. These approaches ignore vital temporal information by simply concatenating features across frames, and thus fail to understand how behavior is changing over time. In contrast, Karácsony et al. [104] employed a 3D-CNN (I3D) and an LSTM to extract spatio-temporal features to distinguish TLE and FLE using IR data. Then, the same authors in [106], extended the architecture to 3 classes including a non-seizure class, and introduced a new pre-processing pipeline, utilizing depth cropping to improve classification performance. Moro et al. [145] extended this architecture to differentiate between non-hyperkinetic seizures and sleep-related paroxysmal events, utilizing a majority voting scheme for the entire video prediction.
@@
Distinguishing epileptic seizures from non-epileptic seizures: A main differential diagnosis of epileptic seizures (ES) is psychogenic non-epileptic seizures (PNES, also known as functional or dissociative seizures, part of the spectrum of functional neurological disorder, FND). PNES have a non-epileptic cause thought to involve psychological and as yet poorly understood neurological mechanisms[26]. Given that both ES and PNES can exhibit rhythmic movements, misdiagnosis is common due to overlapping clinical features observed during these 2 distinct seizure types. It is crucial to differentiate between the two, as their clinical management differs, and misdiagnosis may result in unnecessary treatment and associated complications. We refer here to both diagnostic groups, epileptic seizures (ES) and FND, under the common label of seizure disorders. In the study by Ahmedt-Aristizabal et. al. [4], skeleton-based and spatio-temporal frameworks were compared, with both providing effectiveness in distinguishing between seizure disorders. The authors quantified semiology using either a fusion of reference points and flow fields or by analyzing the entire body simultaneously. While this design is valuable for evaluating dissociative attacks, the region-based approach raises a clinical limitation regarding the observation of isolated semiology. In another work, Hou et al. [87] demonstrated a method to process the appearance and keypoint information for both pose and face streams. They introduced an adaptive GCN where the graph’s topology could be learned on detected body joints and facial landmarks for seizure classification. Knowledge distillation was also adopted to regulate the keypoint features learned by the adaptive GCN. However, including joint information may render the system vulnerable in conditions where joint estimation is poorly performed. In a subsequent work [88], the same authors exploited appearance with a Transformer-based framework. This model was pre-trained on large, unlabeled clinical videos (contextual videos) and then fine-tuned to distinguish between ES and PNES.
@@
Seizure evolution analysis: Recent advances in video analytics have been useful in capturing and quantifying epileptic seizures. However, a critical aspect that has often been overlooked is the representation of the evolving semiology, specifically tracking the stepwise progression of clinical features. Many studies focused on the detection and categorization of movement patterns, neglecting the quantification of multi-segmental rhythmic behaviors and the identification of certain movements, including the order in which they occurred. Understanding the electroclinical patterns of a seizure requires a spatio-temporal profile to be established, allowing for the evaluation of the seizure’s origin and propagation patterns [41]. Approaches incorporating statistical information offer quantitative movement parameters by considering the entire duration of semiology [50]. Previous deep learning frameworks for HAR provide a single result reflecting the classifier’s decision and cannot convey the dynamic changes observed. Ahmedt-Aristizabal et al. [11] developed a system that captures the motion dynamics of body, face and hand semiology, enabling the visualization and analysis of dynamic changes over an entire seizure. This system utilizes a compact image representation known as a motion signature, which captures the spatial location of semiology and the temporal relation between frames. It can also correlate different types of semiology when a patient experiences multiple semiologies simultaneously, thus displaying the order of signals as a stepwise progression. Hou et al. [90] reported that automated video analysis enabled the characterization of rhythmic body movements as a stereotyped expression of frontal seizure semiology. The authors documented time-evolving frequencies as a foundation for studying the correlation between the spectrum of EEG and the frequency of head movements. Developing a system that provides a continuous flow of signals, highlighting changes in semiology and common clinical events, could be a powerful tool to support seizure diagnosis and classification.
@@
Seizure classification (Spatio-temporal-based HAR)
@@
Tasks-constrained deep convolutional network (TCDCN); Position map Regression Network (PRNet); Human Pose Estimation (HPE); Part Affinity Fields (PAF); Convolution Pose Machines (CPM); Long Short Term Memories (LSTM); Feature Engineering (FE) (handcrafted feature extraction); Spatiotemporal interest points (STIP); K-Nearest Neighbors (KNN); Support Vector Machines (SVM); Convolutional Neural Networks (CNNs); Fully Connected layer (FC); Motion Capture (MoCap); Facial emotion recognition (FER); Random Forest (RF); Temporal Segment Networks (TSNs); Spatiotemporal CNN (STCNN); Vision Transformer (ViT); Knowledge Distillation (KD); Progressive Knowledge Distillation (PKD); Adaptive Graph Convolutional Networks (AGCNs); Gated Recurrent Unit (GRU); Multilayer perceptron (MLP); Bidirectional LSTM (Bi-LSTM); $\star$ Human pose detected with depth camera hardware;
Summary: A comprehensive overview of diverse deep learning techniques, their corresponding sources, and the authors from Table 1 who have adapted each technique is presented in Table 2. The related works are systematically categorized according to the frameworks discussed in Section 2.2. The categories include, seizure detection [3, 219, 75, 214, 73, 2, 8, 94, 60, 64, 141, 6], image-based (static) seizure classification [138, 164], skeleton-based seizure classification [50, 10, 5, 9], spatio-temporal seizure classification [10, 5, 160, 4, 227, 104, 105, 106, 145, 162, 88], and hybrid networks for seizure classification [4, 87, 89]. This classification facilitates a structured exploration of the diverse methodologies employed across different aspects of seizure analysis within the deep learning paradigm.
@@ 3.2 General observations regarding vision-based seizure analysis methods
    The automated analysis of video-recorded seizures is a greatly needed to support the detection and identification of epileptic seizures, as well as to better understand the temporal evolution of seizures, from their clinical onset through to termination. While this task has benefited from the use of advanced computer vision and machine learning algorithms, methods are limited by their use of short video clips rather than the entire video of a seizure. This current approach limits the ability of systems to precisely detect the clinical onset, or capture in fine-detail the propagation of semiology.
@@
    For systems to be deployed in a real-world setting it is crucial that practitioners are confident that models can generalize to unseen patients. As such, when developing and evaluating systems researchers must ensure separation of patients between training and testing sets, such that all footage of a given patient lies exclusively in either the training or the testing set to evaluate the generalizability of the proposed system. In some works, this data separation was not considered, which could produce misleadingly high performances. The inclusion of subject-specific features may introduce data leaks, boosting performance without a genuine basis for the target classification.
@@
    To facilitate fair comparisons, recent releases of public datasets based on optical flow representation of different seizure semiologies [162, 141], provide a valuable resource for researchers to conduct standardized evaluations of spatio-temporal HAR approaches. A recent open seizure database that facilitates research into non-EEG seizure detection [163] (accelerometry, photoplethysmography, and pulse-oximetry) is expected to foster collaborative efforts and advance the field of seizure detection. However, there is an increasing need for datasets that provide more comprehensive visual context, particularly regarding behaviors such as facial expressions and hand movements. Striking a balance between data sharing and privacy preservation is challenging yet essential. Possibly, using facial transformation methods in clinical video datasets could mitigate privacy issues and allow datasets of sufficient size to be assembled [86].
@@
    Within the broader deep-learning field there has been a trend towards building end-to-end systems, that is a single system that takes an input and produces the desired output with no intermediate steps or additional outputs. This paradigm is less prevalent within epilepsy analysis, though recent end-to-end transformed based methods concerned with whole body analysis have been proposed. The lack of end-to-end systems likely stems from a combination of the lack of suitably large datasets to train such systems, a reliance on existing methods (i.e. pose detection) to underpin the methods, and also the lack of interpretability that such end-to-end systems offer. It is also worth considering if such end-to-end methods are even appropriate, as processing a whole-body image in an end-to-end fashion is likely to result in hand and face regions are that too small from which to extract fine motor movements.

### 04 Mar 2024 v2

@@ 3.4.2 Action recognition models (frame-based HAR)
Transformer-based methods: As previously noted, 3D CNN approaches have several drawbacks that limit their applicability when dealing with complex actions of variable durations. Transformer-based approaches bridge this gap by providing a unique solution. The Transformer model can process sequential data in parallel by introducing positional embeddings. An early transformer-based method [93], Video Action Transformer Network (VATN), consists of two main components: the Spatial Transformer and the Temporal Transformer. The learned transformers allow the network to selectively attend to important features, resulting in improved action recognition accuracy. Recently, the MTV [249] framework combines multiple views of a video sequence to capture comprehensive spatial and temporal information. This introduces a fusion mechanism that combines the outputs of the transformers. Additionally,  [75] introduces a framework that utilizes masked auto-encoders to capture spatial and temporal dependencies in video data. This approach offers a valuable mechanism for leveraging masked auto-encoders with video data to enhance the understanding and recognition of intricate video patterns. We refer keen readers to [100] and [224] for a more detailed analysis of action recognition methods.
@@ 4.1 Applications and model frameworks
Initial studies have primarily employed hand-designed features based on motion-strength, motion-trajectory, or average differential luminance signals in video segments to develop automated seizure detection systems. For example, hand-crafted features are extracted based on patient motion trajectories by attaching infrared reflective markers to specific body key points and by exploiting RGB-D sensors (e.g. the Microsoft Kinect camera) to distinguish Temporal Lobe Epilepsy (TLE) and Extratemporal Lobe Epilepsy (ETLE) [54]. Aghaei et al. [3] proposed a multimodal seizure detection approach to detect nocturnal convulsive seizures by combining hand-crafted features from video and EEG signals, achieving a higher accuracy than using either modality alone. Similarly, spatio-temporal interest points are extracted from video sequences to construct a histogram of word frequency features for video representation to detect Rolandic Epilepsy in [245]. Other works have focused on nocturnal convulsive seizures using hand-crafted detection algorithms [91, 239] by exploiting optical flow calculations and the reconstruction vector field of velocities [118]. The algorithm quantifies the oscillatory movements seen as vibrations during the tonic phase, and clonic movements in the clonic phase. Recently, Garção et al. [84] proposed a privacy-preserving video-based seizure-detection and registration method to classify generalized tonic–clonic and focal to bilateral tonic–clonic seizures. This approach is based on optical flow, dimensionality reduction with principal component analysis, seizure movement isolation with independent component analysis, and machine learning classification. However, these approaches are limited in performance due to their inability to generalize to changing clinical environments (e.g. changes to subject orientation and distance, dealing with multiple individuals in the field of view, accounting for lighting and furniture changes, motion blur, variable frame rate) and their reliance on handcrafted features.
@@
For early detection, Fang et al. [70] proposed a spatio-temporal gated recurrent unit convolutional neural network capable of not only early detection, but also of localizing the onset time of epileptic seizures. However, the specific seizure types considered and whether the system detects the full expression of the semiology or just the seizure onset remain unclear. In a recent study, Mehta et al. [161] demonstrated the feasibility of detecting tonic-clonic seizures (TCS) during their progression based on optical flow data. The approach utilized knowledge distillation and transformer-based approaches, showcasing a novel direction for early detection while addressing privacy concerns.
@@
The creation of devices capable of objectively recording seizure counts and characteristics holds immense value, not only for seizures occurring during sleep but also for those manifesting while awake. There has been an increasing interest in the analysis of extensive video data collected at home, particularly for understanding seizure semiology in patients using multiple recording modalities [13]. A notable example is the Nelli®hybrid system [23], a semi-automatic seizure monitoring platform utilizing audio/video inputs. This hybrid system employs computer vision and machine learning to identify kinematic data associated with seizures. Human experts then visually assess these identified events. The utility for reviewing nocturnal video recordings and motor seizures has been demonstrated [16, 185, 139]. While Nelli demonstrates proficiency in detecting subtle motor seizures, a study noted a lower accuracy in classifying more discrete motor seizures such as myoclonic jerks, short tonic seizures, and epileptic spasms [178]. The system has received endorsement for clinical use in Finland from a government-established body, the National Coordinating Group for Drug-resistant Epilepsy.
@@
Seizure semiology classification: Understanding the evolution of automated pipelines in the context of video analysis for seizure event classification is crucial. While action recognition requires extracting spatio-temporal features of movements for the classification of epilepsy types, some works overlook the temporal information within seizures, treating the task as an image classification problem. For example, Maia et al. [158] used an InceptionV3 feature extractor to extract features for all frames, which were subsequently concatenated and fed to a multilayer perceptron, to distinguish between TLE and ETLE within IR seizure videos. Similarly, Pothula et al. [184] differentiated between gelastic and dacrystic facial seizures using a feature extraction layer that recognizes the patient’s emotion through a facial emotion recognition model and classification layer (Random Forest) that received concatenated features. These approaches ignore vital temporal information by simply concatenating features across frames, and thus fail to understand how behavior is changing over time (a key aspect for optimal seizure semiology analysis). In contrast, Karácsony et al. [120] employed a 3D-CNN (I3D) and an LSTM to extract spatio-temporal features to distinguish TLE and FLE using IR data. Then, the same authors in [122], extended the architecture to 3 classes including a non-seizure class, and introduced a new pre-processing pipeline, utilizing depth cropping to improve classification performance. Moro et al. [165] extended this architecture to differentiate between non-hyperkinetic seizures and sleep-related paroxysmal events, utilizing a majority voting scheme for the entire video prediction.
@@
Distinguishing epileptic seizures from non-epileptic seizures: A main differential diagnosis of epileptic seizures (ES) is psychogenic non-epileptic seizures (PNES, also known as functional or dissociative seizures, part of the spectrum of functional neurological disorder, FND). PNES have a non-epileptic cause thought to involve psychological and as yet poorly understood neurological mechanisms [27]. Given that both ES and PNES can exhibit rhythmic movements, misdiagnosis is common due to overlapping clinical features observed during these 2 distinct seizure types. It is crucial to differentiate between the two, as their clinical management differs, and misdiagnosis may result in unnecessary treatment and associated complications. We refer in this section to both diagnostic groups, epileptic seizures (ES) and FND, under the common label of seizure disorders. In the study by Ahmedt-Aristizabal et. al. [4], skeleton-based and frame-based frameworks were compared, with both providing effectiveness in distinguishing between seizure disorders. The authors quantified semiology using either a fusion of reference points and flow fields or by analyzing the entire body simultaneously. While this design is valuable for evaluating PNES, the region-based approach raises a clinical limitation regarding the observation of isolated semiology. In another work, Hou et al. [103] demonstrated a method to process the appearance and keypoint information for both pose and face streams. They introduced an adaptive GCN where the graph’s topology could be learned on detected body joints and facial landmarks for seizure classification. Knowledge distillation was also adopted to regulate the keypoint features learned by the adaptive GCN. However, including joint information may render the system vulnerable in conditions where joint estimation is poorly performed. In a subsequent work [104], the same authors exploited appearance with a Transformer-based framework. This model was pre-trained on large, unlabeled clinical videos (contextual videos) and then fine-tuned to distinguish between ES and PNES.
@@
Seizure evolution analysis: Recent advances in video analytics have been useful in capturing and quantifying epileptic seizures. However, a critical aspect that has often been overlooked is the representation of the evolving semiology, specifically tracking the stepwise progression of clinical features. Many studies focused on the detection and categorization of movement patterns, neglecting the quantification of multi-segmental rhythmic behaviors and the identification of certain movements, including the order in which they occurred. Understanding the electroclinical patterns of a seizure requires a spatio-temporal profile to be established, allowing for the evaluation of the seizure’s origin and propagation patterns [45]. Approaches incorporating statistical information offer quantitative movement parameters by considering the entire duration of semiology [54]. Previous deep learning frameworks for HAR provide a single result reflecting the classifier’s decision and cannot convey the dynamic changes observed. Ahmedt-Aristizabal et al. [11] developed a system that captures the motion dynamics of body, face and hand semiology, enabling the visualization and analysis of dynamic changes over an entire seizure. This system utilizes a compact image representation known as a motion signature, which captures the spatial location of semiology and the temporal relation between frames. It can also correlate different types of semiology when a patient experiences multiple semiologies simultaneously, thus displaying the order of signals as a stepwise progression. Hou et al. [106] reported that automated video analysis enabled the characterization of rhythmic body movements as a stereotyped expression of frontal seizure semiology. The authors documented time-evolving frequencies as a foundation for studying the correlation between the spectrum of EEG and the frequency of head movements. Developing a system that provides a continuous flow of signals, highlighting changes in semiology and common clinical events, could be a powerful tool to support seizure diagnosis and classification. Indeed, such developments applied to larger clinical datasets could shed new light on currently unknown neural mechanisms underlying the behavioral expression of seizures.
@@
Seizure classification (Frame-based HAR)
@@
Tasks-constrained deep convolutional network (TCDCN); Position map Regression Network (PRNet); Human Pose Estimation (HPE); Part Affinity Fields (PAF); Convolution Pose Machines (CPM); Long Short Term Memories (LSTM); Feature Engineering (FE) (handcrafted feature extraction); Spatio-temporal interest points (STIP); K-Nearest Neighbors (KNN); Support Vector Machines (SVM); Convolutional Neural Networks (CNNs); Fully Connected layer (FC); Motion Capture (MoCap); Facial emotion recognition (FER); Random Forest (RF); Temporal Segment Networks (TSNs); Spatio-temporal CNN (STCNN); Vision Transformer (ViT); Knowledge Distillation (KD); Progressive Knowledge Distillation (PKD); Adaptive Graph Convolutional Networks (AGCNs); Gated Recurrent Unit (GRU); Multilayer perceptron (MLP); Bidirectional LSTM (Bi-LSTM); $\star$ Human pose detected with depth camera hardware;
Summary: A comprehensive overview of diverse deep learning techniques, their corresponding sources, and the authors from Table 1 who have adapted each technique is presented in Table 2. The related works are systematically categorized according to the frameworks discussed in Section 3.2. The categories include, seizure detection [3, 245, 91, 239, 84, 2, 8, 110, 66, 70, 161, 6], image-based (static) seizure classification [158, 184], skeleton-based seizure classification [54, 10, 5, 9], frame-based seizure classification [10, 5, 180, 4, 253, 120, 121, 122, 165, 182, 104], and hybrid networks for seizure classification [4, 103, 105]. This classification facilitates a structured exploration of the diverse methodologies employed across different aspects of seizure analysis within the deep learning paradigm.
@@ 4.2 General observations regarding vision-based seizure analysis methods
    The automated analysis of video-recorded seizures is greatly needed to support the detection and identification of epileptic seizures, as well as to better understand the temporal evolution of seizures, from their clinical onset through to termination. While this task has benefited from the use of advanced computer vision and machine learning algorithms, methods are limited by their use of short video clips rather than the entire video of a seizure. This current approach limits the ability of systems to precisely detect the clinical onset, or capture in fine-detail the propagation of semiology.
@@
    For systems to be deployed in a real-world setting it is crucial that practitioners are confident that models can generalize to unseen patients. This is particularly important for analyzing complex motor behaviors in seizures, since the overall behavioral repertoire and inter-subject and even intra-subject variability is considerable. As such, when developing and evaluating systems researchers must ensure separation of patients between training and testing sets, such that all footage of a given patient lies exclusively in either the training or the testing set to evaluate the generalizability of the proposed system. In some works, this data separation was not considered, which could produce misleadingly high performances. The inclusion of subject-specific features may introduce data leaks, boosting performance without a genuine basis for the target classification.
@@
    To facilitate fair comparisons, recent releases of public datasets based on optical flow representation of different seizure semiologies [182, 161], provide a valuable resource for researchers to conduct standardized evaluations of frame-based HAR approaches. A recent open seizure database that facilitates research into non-EEG seizure detection [183] (accelerometry, photoplethysmography, and pulse-oximetry) is expected to foster collaborative efforts and advance the field of seizure detection. However, there is an increasing need for datasets that provide more comprehensive visual context, particularly regarding behaviors such as facial expressions and hand movements. Striking a balance between data sharing and privacy preservation is challenging yet essential. Possibly, using facial transformation methods in clinical video datasets could mitigate privacy issues and allow datasets of sufficient size to be assembled [102].
@@
    Within the broader deep-learning field there has been a trend towards building end-to-end systems, that is a single system that takes an input and produces the desired output with no intermediate steps or additional outputs. This paradigm is less prevalent within Epilepsy analysis, though recent end-to-end transformer-based methods concerned with whole body analysis have been proposed. The lack of end-to-end systems likely stems from a combination of the lack of suitably large datasets to train such systems, a reliance on existing methods (i.e. pose detection) to underpin the methods, and also the lack of interpretability that such end-to-end systems offer. It is also worth considering if such end-to-end methods are even appropriate, as processing a whole-body image in an end-to-end fashion is likely to result in hand and face regions are that too small from which to extract fine motor movements.